Solve using the five-step problem-solving process.
The sum of two consecutive integers is -247. Find the larger integer.

Answer by checkley77(12844)

Let x & x+1 be the 2 integers.
x+(x+1)=-247
2x+1=-247
2x=-247-1
2x=-248
x=-248/2
x=-124 ans. for the smaller integer.
-124+1=-123 ans. for the larger integer.
Proof:
-124-123=-247